# Precompactness in matrix weighted Bourgain-Morrey spaces

### v1

@@
Abstract.
In this paper, we introduce matrix weighted Bourgain-Morrey spaces and obtain two sufficient conditions for precompact sets in matrix weighted Bourgain-Morrey spaces. We prove that the dyadic average operator is bounded on some matrix weighted Bourgain-Morrey spaces. With this result, we obtain the necessity for precompact sets in some matrix weighted Bourgain-Morrey spaces.
@@ 1 Introduction
The theory of matrix weighted function spaces goes back to [32]. Indeed, in 1958, Wiener and Masani [32, Section 4] studied the matrix weighted $L^{2}(W)$ for the prediction theory for multivariate stochastic processes. In [29], Treil and Volberg introduced matrix class $\mathcal{A}_{2}$. Nazarov and Treil [23] and Volberg [30] extended $\mathcal{A}_{2}$ to $\mathcal{A}_{p}$ with $p\in(1,\infty)$. In [11], Goldberg showed that the matrix $\mathcal{A}_{p}$ condition leads to $L^{p}$ boundedness of a Hardy-Littlewood maximal function and obtained the boundedness of matrix weighted singular integral operators in Lebesgue spaces $L^{p},1<p<\infty$. In [24, 25, 26], Roudenko introduced the matrix-weighted homogeneous Besov spaces $\dot{B}^{s,q}_{p}(W)$ and matrix-weighted sequence Besov spaces $\dot{b}^{s,q}_{p}(W)$ and showed their equivalence via $\varphi$-transform and wavelets. In [9], Frazier and Roudenko introduced the matrix class $\mathcal{A}_{p},(0<p\leq 1)$ and studied the continuous and discrete matrix-weighted Besov spaces $\dot{B}^{s,q}_{p}(W)$ and $\dot{b}^{s,q}_{p}(W)$ with $0<p\leq 1$. In [10], Frazier and Roudenko introduced the homogeneous matrix-weighted Triebel-Lizorkin spaces $\dot{F}^{s,q}_{p}(W)$ for $s\in\mathbb{R},0<p<\infty,0<q\leq\infty$ and obtained the Littlewood-Paley characterizations of matrix-weighted Lebesgue spaces $L^{p}(W)$ and matrix-weighted Sobolev space $L^{p}_{k}(W)$ for $k\in\mathbb{N},1<p<\infty$. In [30], Volberg also introduced an analogue condition for the matrix class $\mathcal{A}_{\infty}$. In [3, 4, 5], Bu, Hytönen, Yang, Yuan studied the matrix weighted Besov-type and Triebel-Lizorkin-type spaces. Specifically, they introduced a new concept of the $\mathcal{A}_{p}$-dimension $\tilde{d}$, which is useful to the proof of main results of this paper. In [2], Bu, Hytönen, Yang, Yuan obtained several new characterizations of $\mathcal{A}_{p,\infty}$-matrix weights. For many other results on the matrix class $\mathcal{A}_{p}$ and matrix weighted function spaces, we refer the reader to [6, 8, 9, 10, 11, 24, 25, 31].
@@
Motivated by above literature, we will introduce matrix weighted Bourgain-Morrey spaces and research precompact sets in these spaces. The paper is organized as follows. In Section 2, dyadic cubes, the matrix class $\mathcal{A}_{p}$, $\mathcal{A}_{p}$-dimension $\tilde{d}$, matrix weighted Bourgain-Morrey spaces are given. The first result lies in Section 3. Specifically, a sufficient condition for totally bounded set in matrix weighted spaces $M_{p}^{t,r}(W)$ with $1\leq p<r<\infty$ or $1\leq p\leq t<r=\infty$ is obtained in Theorem 3.2. As a application, we get a criterion for totally bounded set in degenerate Bourgain-Morrey spaces with matrix weight. The second result (Theorem 4.1) is replacing the translation operator by the average operator. Note that the translation operator is not bounded on $L^{p}(W)$ and $M_{p}^{t,r}(W)$ in general. We prove that dyadic average operator is bounded on matrix weighted Bourgain-Morrey spaces with some conditions in Theorem 4.5. Using this result, we obtain the Kolmogorov-Riesz compactness theorem in matrix weighted Bourgain-Morrey spaces.
Throughout this paper, we let $c,C$ denote constants that are independent of the main parameters involved but whose value may differ from line to line. Let =$\mathbb{N}=\{1,2,3,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\ldots\}$. Let $\mathbb{Z}$ be the set of integers. Let $\chi_{E}$ be the characteristic function of the set $E\subset{\mathbb{R}^{n}}$. By $A\lesssim B$ we mean that $A\leq CB$ with some positive constant $C$ independent of appropriate quantities. By $A\approx B$, we mean that $A\lesssim B$ and $B\lesssim A$.
@@ 2 Preliminaries
For $j\in\mathbb{Z}$, $m\in\mathbb{Z}^{n}$, let $Q_{j,m}:=\prod_{i=1}^{n}[2^{-j}m_{i},2^{-j}(m_{i}+1))$. For a cube $Q$, $\ell(Q)$ stands for the length of cube $Q$. We denote by $\mathcal{D}$ the the family of all dyadic cubes in $\mathbb{R}^{n}$, while $\mathcal{D}_{j}$ is the set of all dyadic cubes with $\ell(Q)=2^{-j},j\in\mathbb{Z}$. Let $x_{Q}$ be the left corner of $Q\in{\mathcal{D}}$. For $\lambda>0$, let $\lambda Q$ be the cube with the same center of $Q$ and the edge length $\lambda\ell(Q)$. For $k\in\mathbb{N}$, let $k_{\operatorname{pa}}Q$ be the $k$-th dyadic parent of $Q$, which is the dyadic cube in ${\mathcal{D}}$ satisfying $Q\subset k_{\operatorname{pa}}Q$ and $\ell(k_{\operatorname{pa}}Q)=2^{k}\ell(Q)$.
@@ 2.1 Matrix weights
We say that a matrix $A\in M_{d}(\mathbb{C})$ is positive definite if, for any $\vec{x}\in\mathbb{C}^{d}\backslash\{\vec{0}\}$, $\vec{z}^{*}A\vec{z}>0.$ And a matrix $A\in M_{d}(\mathbb{C})$ is called nonnegative definite if, for any $\vec{x}\in\mathbb{C}^{d}$, $\vec{z}^{*}A\vec{z}\geq 0.$
From [18, Theorem 4.1.4], any nonnegative definite matrix is always Hermitian. Hence any nonnegative definite matrix is self-adjoint.
Let $A\in M_{d}(\mathbb{C})$ be a positive definite matrix and have eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$. From [18, Theorem 2.5.6(c)], there exists a unitary matrix $U\in M_{d}(\mathbb{C})$ such that
$A=U\operatorname{diag}\,(\lambda_{1},\ldots,\lambda_{d})U^{*}.$ (2.2)
Moreover, by [18, Theorem 4.1.8], we find $\{\lambda_{i}\}_{i=1}^{d}\subset(0,\infty)^{d}$. The following definition is based on these conclusions.
Definition 2.1
@@
$A^{\alpha}:=U\operatorname{diag}\left(\lambda_{1}^{\alpha},\ldots,\lambda_{m}^{\alpha}\right)U^{*},$ (2.3)
where $U$ is the same as in (2.2).
Remark 2.2
From [17, p. 408], we obtain that $A^{\alpha}$ is independent of the choices of the order of $\{\lambda_{i}\}_{i=1}^{m}$ and $U$, and hence $A^{\alpha}$ is well defined.
@@
(iii) the entries of $W$ is locally integrable;
Definition 2.4
@@
$\left|A_{E}\vec{z}\right|\approx\left(\frac{1}{|E|}\int_{E}\left|W^{\frac{1}{p}}(x)\vec{z}\right|^{p}\,{\mathrm{d}}x\right)^{\frac{1}{p}},$ (2.4)
@@
Next we recall the concepts of scalar weight class $A_{p}$ (see [12, Definitions 7.1.1, 7.1.3]) and matrix weight class $\mathcal{A}_{p}$ (see [26] for $1<p<\infty$, [9] for $0<p\leq 1$).
Definition 2.5
@@
[…]
@@
Given any matrix weight $W$ and $0<p<\infty$, there exists (see e.g., [11, Proposition 1.2]) for $p>1$ and [9, p.1237] for $0<p\leq 1)$ a sequence $\{A_{Q}\}_{Q\in\mathcal{D}}$ of positive definite $d\times d$ matrices such that
@@
From [15, Proposition 2.10], we know that $\beta$ is always not less than $n$.
In [3], Bu, Hytönen, Yang and Yuan introduced the $\mathcal{A}_{p}$-dimension of matrix weights, which will be used in Theorems 4.5 and 4.6.
Definition 2.7
@@
[…]
@@
[…]
where $11/p+1/p^{\prime}=$ 1.
@@
Lemma 2.8 (Corollary 2.32, [3])
Let $0<p<\infty$, let $W\in\mathcal{A}_{p}$ have the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$, and let $\{A_{Q}\}_{\operatorname{cube}Q}$ be the reducing operator of order $p$ for $W$. (i) If $1<p<\infty$, let $\widetilde{W}:=W^{-1/(p-1)}$ (which belongs to $\mathcal{A}_{p^{\prime}}$) have the $\mathcal{A}_{p^{\prime}}$-dimension $\tilde{\tilde{d}}$. Then there exists a positive constant $C$ such that, for any cubes $Q$ and $R$ of ${\mathbb{R}^{n}}$,
@@ 2.2 Matrix weighted Bourgain-Morrey spaces
Let $\varOmega\subset{\mathbb{R}^{n}}$ be an open set. The space $L^{p}(W,\varOmega)$ collects all measurable functions $\vec{f}:\varOmega\to\mathbb{C}^{d}$ such that
$\|\vec{f}\|_{L^{p}(W,\varOmega)}=\left(\int_{\varOmega}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}<\infty.$
Definition 2.10
@@
Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$. Define $M_{p}^{t,r}(\{A_{Q}\})$ as the set of all $\vec{f}\in L_{\rm loc}^{p}(\{W\})$ such that
$\|\vec{f}\|_{M_{p}^{t,r}(\{A_{Q}\})}$
:=$:=\left\|\left\{\left(\|A_{Q_{j,k}}\|^{p}|Q_{j,k}|\right)^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(y)|^{p}{\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$
Remark 2.11
Let $\{\vec{e}_{i}\}_{i=1}^{d}$ be the standard unit basis for $\mathbb{C}^{d}$. Since $W$ is a matrix weight, then for each $x\in{\mathbb{R}^{n}}$, $W(x)$ is a nonnegative matrix and self-adjoint. Hence we have
$\|W(x)\|$ $=\|W^{2/p}(x)\|^{p/2}\approx[\operatorname{tr}(W^{2/p}(x))]^{p/2}$
$=\left(\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{2}\right)^{p/2}\approx\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{p}.$
From this and (2.4), we obtain
$W(Q_{j,k})$ $=\int_{Q_{j,k}}\|W(y)\|{\mathrm{d}}y\approx\int_{Q_{j,k}}\sum_{i=1}^{d}|W^{1/p}(x)\vec{e}_{i}|^{p}{\mathrm{d}}y$
$\approx\sum_{i=1}^{d}|A_{Q_{j,k}}\vec{e}_{i}|^{p}|Q_{j,k}|\approx\|A_{Q_{j,k}}\|^{p}|Q_{j,k}|.$
@@
$\|f\|_{M_{p}^{t,r}(\omega)}:=\left\|\left\{(\omega(Q_{j,k}))^{1/t-1/p}\left(\int_{Q_{j,k}}|f(y)|^{p}\omega(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$ (2.6)
@@
Lemma 2.12
Let $1\leq p<\infty$ and $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. Let $\varOmega\subset{\mathbb{R}^{n}}$ be an open set. Then matrix weighted Lebesgue space $L^{p}(W,\varOmega)$ is a Banach space.
Theorem 2.13
@@
  • Proof
    Let $\{\vec{f}_{k}\}_{k=1}^{\infty}$ be a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$. That is, for any compact set $K$, any $\varepsilon>0$, there exists $N>0$ such that if $j,k>N$, $\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\varepsilon.$ Then for any compact set $K$, $\{\vec{f}_{k}\chi_{K}\}_{k=1}^{\infty}$ is a Cauchy sequence in $L^{p}(W)$. Since $L^{p}(W)$ is complete (Lemma 2.12), there exists $\vec{f}\chi_{K}$ in $L^{p}(W)$ such that $\vec{f}_{k}\chi_{K}\to\vec{f}\chi_{K}$. By the Fatou lemma, we have
@@
    $\|(\vec{f}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}=\lim_{j\to\infty}\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\varepsilon.$
    Hence $\vec{f}_{k}\to\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ and the space $L^{p}_{\operatorname{loc}}(W)$ is complete. $\Box$
@@
    $\leq\left(\frac{1}{W(Q)}\int_{Q}\|W^{1/p_{1}-1/p_{2}}(y)\|^{p_{1}p_{2}/(p_{2}-p_{1})}{\mathrm{d}}y\right)^{1/p_{1}-1/p_{2}}$
    ×$\quad\times\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{2}}{\mathrm{d}}y\right)^{1/p_{2}}$
@@
    For any compact set $K\subset{\mathbb{R}^{n}}$, there exist at most $2^{n}$ dyadic cubes $Q_{j}$ such that $K\subset\cup_{j=1}^{2n}Q_{j}$. Hence
    $\|\vec{f}\chi_{K}\|_{L^{p}(W)}$ $\leq\sum_{j=1}^{2^{n}}\frac{|Q_{j}|^{1/t-1/p}}{|Q_{j}|^{1/t-1/p}}\|\vec{f}\chi_{Q_{j}}\|_{L^{p}(W)}$
    $\leq\|\vec{f}\|_{M_{P}^{t,\infty}(W)}\sum_{j=1}^{2^{n}}\frac{1}{|Q_{j}|^{1/t-1/p}}<\infty.$
@@
  • Proof
    The argument are standard, see, for example, [27, Theorem 2.4]. We only show that $M_{p}^{t,r}(W)$ is complete since others are simple.
Let $\{\vec{f}_{j}\}_{j=1}^{\infty}$ be a Cauchy sequence in $M_{p}^{t,r}(W)$. By Proposition 1, we have that $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is also a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$. Since $L^{p}_{\operatorname{loc}}(W)$ is complete, there exists a vector function $\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ such that $\vec{f}_{j}\to\vec{f}$. By the Fatou lemma,
@@
Thus, the sequence $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is convergent to $\vec{f}$ in $M_{p}^{t,r}(W)$. $\Box$

### v2

@@
Abstract
In this paper, we introduce matrix weighted Bourgain-Morrey spaces and obtain two sufficient conditions for precompact sets in matrix weighted Bourgain-Morrey spaces. We prove that the dyadic average operator is bounded on some matrix weighted Bourgain-Morrey spaces. With this result, we obtain the necessity for precompact sets in some matrix weighted Bourgain-Morrey spaces. The results are new even for the unweighted Bourgain-Morrey spaces.
Keywords Kolmogorov-Riesz theorem, matrix weight, precompactness, Bourgain-Morrey space.
Mathematics Subject Classification Primary 46B50, 46E40, 42B35, 46E30.
@@ 1 Introduction
The theory of matrix weighted function spaces goes back to [35]. Indeed, in 1958, Wiener and Masani [35, Section 4] studied the matrix weighted $L^{2}(W)$ for the prediction theory for multivariate stochastic processes. In [32], Treil and Volberg introduced matrix class $\mathcal{A}_{2}$. Nazarov and Treil [26] and Volberg [33] extended $\mathcal{A}_{2}$ to $\mathcal{A}_{p}$ with $p\in(1,\infty)$. In [13], Goldberg showed that the matrix $\mathcal{A}_{p}$ condition leads to $L^{p}$ boundedness of a Hardy-Littlewood maximal function and obtained the boundedness of matrix weighted singular integral operators in Lebesgue spaces $L^{p},1<p<\infty$. In [28, 29, 27], Roudenko introduced the matrix-weighted homogeneous Besov spaces $\dot{B}^{s,q}_{p}(W)$ and matrix-weighted sequence Besov spaces $\dot{b}^{s,q}_{p}(W)$ and showed their equivalence via $\varphi$-transform and wavelets. In [11], Frazier and Roudenko introduced the matrix class $\mathcal{A}_{p},(0<p\leq 1)$ and studied the continuous and discrete matrix-weighted Besov spaces $\dot{B}^{s,q}_{p}(W)$ and $\dot{b}^{s,q}_{p}(W)$ with $0<p\leq 1$. In [12], Frazier and Roudenko introduced the homogeneous matrix-weighted Triebel-Lizorkin spaces $\dot{F}^{s,q}_{p}(W)$ for $s\in\mathbb{R},0<p<\infty,0<q\leq\infty$ and obtained the Littlewood-Paley characterizations of matrix-weighted Lebesgue spaces $L^{p}(W)$ and matrix-weighted Sobolev space $L^{p}_{k}(W)$ for $k\in\mathbb{N},1<p<\infty$. In [33], Volberg also introduced an analogue condition for the matrix class $\mathcal{A}_{\infty}$. In [3, 4, 5], Bu, Hytönen, Yang, Yuan studied the matrix weighted Besov-type and Triebel-Lizorkin-type spaces. Specifically, they introduced a new concept of the $\mathcal{A}_{p}$-dimension $\tilde{d}$, which is useful to the proof of main results of this paper. In [7], Bu, Hytönen, Yang, Yuan obtained several new characterizations of $\mathcal{A}_{p,\infty}$-matrix weights. In [8], Bu et al. studied the inhomogeneous Besov-type and Triebel–Lizorkin-type spaces with the result in [7]. In [2], Bu et al. introduced the matrix weighted Hardy spaces and obtained characterizations of these spaces via maximal function, atom. As applications, they established the finite atomic characterization of matrix weighted Hardy spaces and obtained a criterion on the boundedness of sublinear operators from matrix weighted Hardy spaces to any $\gamma$-quasi-Banach space ($\gamma\in(0,1]$). The boundedness of Calderón-Zygmund operators on matrix weighted Hardy spaces was also obtained. In [23], Li, Yang and Yuan introduced the matrix-weighted Besov-Triebel-Lizorkin spaces with logarithmic smoothness and characterize these spaces via Peetre-type maximal functions. In [37, 38], Zhao et al. introduced (generalized grand) Besov-Bourgain-Morrey spaces and explored various real-variable properties of these spaces, which are a bridge connecting Bourgain-Morrey spaces with amalgam-type spaces. Moreover, some real-variable properties and boundedness of classical operators were studied in their article. For many other results on the matrix class $\mathcal{A}_{p}$ and matrix weighted function spaces, we refer the reader to [6, 10, 11, 12, 13, 28, 29, 34].
@@
Motivated by above literature, we will introduce matrix weighted Bourgain-Morrey spaces and research precompact sets in these spaces. The paper is organized as follows. In Section 2, dyadic cubes, the matrix class $\mathcal{A}_{p}$, $\mathcal{A}_{p}$-dimension $\tilde{d}$, matrix weighted Bourgain-Morrey spaces are given. The first result lies in Section 3. Specifically, a sufficient condition for totally bounded set in matrix weighted spaces $M_{p}^{t,r}(W)$ with $1\leq p<r<\infty$ or $1\leq p\leq t<r=\infty$ is obtained in Theorem 3.2. As a application, we get a criterion for totally bounded set in degenerate Bourgain-Morrey spaces with matrix weight. The second result (Theorem 4.1) is replacing the translation operator by the average operator. Note that the translation operator is not bounded on $L^{p}(W)$ and $M_{p}^{t,r}(W)$ in general. We prove that dyadic average operator is bounded on matrix weighted Bourgain-Morrey spaces with some conditions in Theorem 4.6. Using this result, we obtain the Kolmogorov-Riesz compactness theorem in matrix weighted Bourgain-Morrey spaces. These results are new even for the unweighted Bourgain-Morrey spaces.
Throughout this paper, we let $c,C$ denote constants that are independent of the main parameters involved but whose value may differ from line to line. Let $\mathbb{N}=\{1,2,3,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,3,\ldots\}$. Let $\mathbb{Z}$ be the set of integers. Let $\chi_{E}$ be the characteristic function of the set $E\subset{\mathbb{R}^{n}}$. By $A\lesssim B$ we mean that $A\leq CB$ with some positive constant $C$ independent of appropriate quantities. By $A\approx B$, we mean that $A\lesssim B$ and $B\lesssim A$.
@@ 2 Preliminaries
For $j\in\mathbb{Z}$, $m\in\mathbb{Z}^{n}$, let $Q_{j,m}:=\prod_{i=1}^{n}[2^{-j}m_{i},2^{-j}(m_{i}+1))$. For a cube $Q$, $\ell(Q)$ stands for the length of cube $Q$. We denote by $\mathcal{D}$ the the family of all dyadic cubes in $\mathbb{R}^{n}$, while $\mathcal{D}_{j}$ is the set of all dyadic cubes with $\ell(Q)=2^{-j},j\in\mathbb{Z}$. Let $x_{Q}$ be the lower left corner of $Q\in{\mathcal{D}}$. For $\lambda>0$, let $\lambda Q$ be the cube with the same center of $Q$ and the edge length $\lambda\ell(Q)$. For $k\in\mathbb{N}$, let $k_{\operatorname{pa}}Q$ be the $k$-th dyadic parent of $Q$, which is the dyadic cube in $\mathcal{D}$} satisfying $Q\subset k_{\operatorname{pa}}Q$ and $\ell(k_{\operatorname{pa}}Q)=2^{k}\ell(Q)$.
@@ 2.1 Matrix weights
[…]
From [20, Theorem 4.1.4], any nonnegative definite matrix is always Hermitian. Hence any nonnegative definite matrix is self-adjoint.
Let $A\in M_{d}(\mathbb{C})$ be a positive definite matrix and have eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$. From [20, Theorem 2.5.6(c)], there exists a unitary matrix $U\in M_{d}(\mathbb{C})$ such that
$A=U\operatorname{diag}\,(\lambda_{1},\ldots,\lambda_{d})U^{*}.$ (2)
Moreover, by [20, Theorem 4.1.8], we find $\{\lambda_{i}\}_{i=1}^{d}\subset(0,\infty)$. The following definition is based on these conclusions.
Definition 2.1.
@@
$A^{\alpha}:=U\operatorname{diag}\left(\lambda_{1}^{\alpha},\ldots,\lambda_{m}^{\alpha}\right)U^{*},$ (3)
where $U$ is the same as in (2).
Remark 2.2.
From [19, p. 408], we obtain that $A^{\alpha}$ is independent of the choices of the order of $\{\lambda_{i}\}_{i=1}^{m}$ and $U$, and hence $A^{\alpha}$ is well defined.
@@
(iii) the entries of $W$ is locally integrable.
Definition 2.4.
@@
$\left|A_{E}\vec{z}\right|\approx\left(\frac{1}{|E|}\int_{E}\left|W^{\frac{1}{p}}(x)\vec{z}\right|^{p}\,{\mathrm{d}}x\right)^{\frac{1}{p}},$ (4)
@@
Next we recall the concepts of scalar weight class $A_{p}$ (see [14, Definitions 7.1.1, 7.1.3]) and matrix weight class $\mathcal{A}_{p}$ (see [27] for $p<\infty, [11$] for $0<p\leq 1$).
Definition 2.5.
@@
$\underset{Q}{\sup}\;\underset{y\in Q}{\mathop{\mathrm{ess\;sup}}}\frac{1}{|Q|}\int_{Q}\|W^{1/p}(x)W^{-1/p}(y)\|^{p}\mathrm{d}x<\infty.$
@@
Given any matrix weight $W$ and $0<p<\infty$, there exists (see e.g., [13, Proposition 1.2] for $p>1$ and [11, p.1237] for $0<p\leq 1)$ a sequence $\{A_{Q}\}_{Q\in\mathcal{D}}$ of positive definite $d\times d$ matrices such that
@@
From [17, Proposition 2.10], we know that $\beta$ is always not less than $n$.
In [3], Bu, Hytönen, Yang and Yuan introduced the $\mathcal{A}_{p}$-dimension of matrix weights, which will be used in Theorems 4.6 and 4.7.
Definition 2.7.
@@
$\underset{y\in 2^{i}Q}{\operatorname{ess\;\sup}}\frac{1}{|Q|}\int_{Q}\left\|W^{1/p}(x)W^{-1/p}(y)\right\|^{p}{\mathrm{d}}x\leq C2^{i\tilde{d}},\quad\operatorname{for}\;0<p\leq 1,$ (5)
@@
$\frac{1}{|Q|}\int_{Q}\left(\frac{1}{|2^{i}Q|}\int_{2^{i}Q}\left\|W^{1/p}(x)W^{-1/p}(y)\right\|^{p^{\prime}}{\mathrm{d}}y\right)^{p/p^{\prime}}{\mathrm{d}}x\leq C2^{i\tilde{d}},\quad\operatorname{for}\;1<p<\infty,$
where $1/p+1/p^{\prime}=1$.
@@
The following lemma says that if $W\in\mathcal{A}_{p}$ for $p\in(0,\infty)$, then $W\in\mathbb{D}_{p,\tilde{d}}(\mathbb{R}^{n},\mathbb{C}^{n})$.
Lemma 2.8 (Proposition 2.27, [3]).
Let $p\in(0,\infty)$ and $W\in\mathcal{A}_{p}$. Then there exists $\tilde{d}\in[0,n)$ such that $W$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}$.
Lemma 2.9 (Corollary 2.32, [3]).
Let $0<p<\infty$, let $W\in\mathcal{A}_{p}$ with the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$, and let $\{A_{Q}\}_{\operatorname{cube}Q}$ be the reducing operator of order $p$ for $W$. (i) If $1<p<\infty$, let $\widetilde{W}:=W^{-1/(p-1)}$ (which belongs to $\mathcal{A}_{p^{\prime}}$) with the $\mathcal{A}_{p^{\prime}}$-dimension $\tilde{\tilde{d}}$. Then there exists a positive constant $C$ such that, for any cubes $Q$ and $R$ of $\mathbb{R}^{n}$},
@@ 2.2 Matrix weighted Bourgain-Morrey spaces
Let $\Omega\subset{\mathbb{R}^{n}}$ be an open set. The space $L^{p}(W,\Omega)$ collects all measurable functions $\vec{f}:\Omega\to\mathbb{C}^{d}$ such that
$\|\vec{f}\|_{L^{p}(W,\Omega)}=\left(\int_{\Omega}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}<\infty.$
Definition 2.11.
@@
Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$. Define $M_{p}^{t,r}(\{A_{Q}\})$ as the set of all $\vec{f}\in L_{\rm loc}^{p}(W)$ such that
$\|\vec{f}\|_{M_{p}^{t,r}(\{A_{Q}\})}:=\left\|\left\{\left(\|A_{Q_{j,k}}\|^{p}|Q_{j,k}|\right)^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(y)|^{p}{\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$
Remark 2.12.
By [3, Lemma 2.11] with $M=I_{m}$, we conclude that, for any cube $Q\subset{\mathbb{R}^{n}}$,
$\|A_{Q}\|^{p}\approx\frac{1}{|Q|}\int_{Q}\left\|W^{1/p}(x)\right\|^{p}{\mathrm{d}}x=\frac{1}{|Q|}\int_{Q}\left\|W(x)\right\|{\mathrm{d}}x=\frac{1}{|Q|}W(Q).$
@@
$\|f\|_{M_{p}^{t,r}(\omega)}:=\left\|\left\{(\omega(Q_{j,k}))^{1/t-1/p}\left(\int_{Q_{j,k}}|f(y)|^{p}\omega(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}<\infty.$ (6)
@@
Lemma 2.13.
Let $1\leq p<\infty$ and $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$ be a matrix weight. Let $\Omega\subset{\mathbb{R}^{n}}$ be an open set. Then matrix weighted Lebesgue space $L^{p}(W,\Omega)$ is a Banach space.
Theorem 2.14.
@@
Proof.
Let $\{\vec{f}_{k}\}_{k=1}^{\infty}$ be a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$. That is, for any compact set $K$, any $\epsilon>0$, there exists $N>0$ such that if $j,k>N$, $\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\epsilon.$ Then for any compact set $K$, $\{\vec{f}_{k}\chi_{K}\}_{k=1}^{\infty}$ is a Cauchy sequence in $L^{p}(W)$. Since $L^{p}(W)$ is complete (Lemma 2.13), there exists $\vec{f}\chi_{K}$ in $L^{p}(W)$ such that $\vec{f}_{k}\chi_{K}\to\vec{f}\chi_{K}$. By the Fatou lemma, we have
@@
$\|(\vec{f}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}=\lim_{j\to\infty}\|(\vec{f}_{j}-\vec{f}_{k})\chi_{K}\|_{L^{p}(W)}<\epsilon.$
Hence $\vec{f}_{k}\to\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ and the space $L^{p}_{\operatorname{loc}}(W)$ is complete. ∎
@@
$\leq\left(\frac{1}{W(Q)}\int_{Q}\|W^{1/p_{1}-1/p_{2}}(y)\|^{p_{1}p_{2}/(p_{2}-p_{1})}{\mathrm{d}}y\right)^{1/p_{1}-1/p_{2}}\left(\frac{1}{W(Q)}\int_{Q}|W^{1/p_{2}}(y)\vec{f}(y)|^{p_{2}}{\mathrm{d}}y\right)^{1/p_{2}}$
@@
For any compact set $K\subset{\mathbb{R}^{n}}$, there exist at most $2^{n}$ dyadic cubes $Q_{j}$ such that $K\subset\bigcup_{j=1}^{2^{n}}Q_{j}$. Hence
$\|\vec{f}\chi_{K}\|_{L^{p}(W)}\leq\sum_{j=1}^{2^{n}}\frac{|Q_{j}|^{1/t-1/p}}{|Q_{j}|^{1/t-1/p}}\|\vec{f}\chi_{Q_{j}}\|_{L^{p}(W)}\leq\|\vec{f}\|_{M_{P}^{t,\infty}(W)}\sum_{j=1}^{2^{n}}\frac{1}{|Q_{j}|^{1/t-1/p}}<\infty.$
@@
Proof.
The argument are standard, see, for example, [30, Theorem 2.4]. We only show that $M_{p}^{t,r}(W)$ is complete since others are simple.
Let $\{\vec{f}_{j}\}_{j=1}^{\infty}$ be a Cauchy sequence in $M_{p}^{t,r}(W)$. By Proposition 2.15, we have that $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is also a Cauchy sequence in $L^{p}_{\operatorname{loc}}(W)$. Since $L^{p}_{\operatorname{loc}}(W)$ is complete, there exists a vector function $\vec{f}$ in $L^{p}_{\operatorname{loc}}(W)$ such that $\vec{f}_{j}\to\vec{f}$. By the Fatou lemma,
@@
Thus, the sequence $\{\vec{f}_{j}\}_{j=1}^{\infty}$ is convergent to $\vec{f}$ in $M_{p}^{t,r}(W)$. ∎